4. Nitrogen with a pressure of 1.0×105Pa and a volume of 0.0082m3 from the initial temperature of 300K. This gas is heated to 400K according to the following two processes. How much heat does it need to absorb? Which process requires more heat? why?

 (1) The volume does not change, (2) The pressure does not change


1
Expert's answer
2021-07-05T17:59:28-0400

We know that, ΔH = ΔU + PΔV (+ VΔP, ΔP=0 for constant pressure) . So the enthalpy term is greater than the internal energy term because of the PΔV term i.e in case of a constant pressure process more energy is needed, to be provided to the system as compared to that of a constant volume process to achieve the same temperature rise, as some energy is utilized in the expansion work of the system. And the relation that correlates these two is Cp = Cv + R.


1)


"nR=\frac{0.0082(10^5)}{300}=2.7"

"Q=3.5(2.7)(400-300)=945\ J"

2)


"Q=2.5(2.7)(400-300)=675\ J"